OpenStudy (anonymous):

f(x)=x^4-14x^3+72x^2+9x-10 find the largest interval on which f(x) is concave downward. If we write the interval as (a,b), find a and b

OpenStudy (amistre64):

in other words, find the 2 inflection points

OpenStudy (amistre64):

set f'' to 0 and that should do it for this ....

OpenStudy (anonymous):

okay can you just tell me how to do it. i dont understand you

OpenStudy (amistre64):

if the site doesnt die on me i might be able to scrounge up someting :)

OpenStudy (amistre64):

\[\frac{d^2}{dx^2}x^n\to n(n-1)x^{(n-2)}\] is all you are doing to this

OpenStudy (amistre64):

\[f(x)=x^4-14x^3+72x^2+9x-10\] \[f''(x)=4.3x^2-14.3.2x+72.2.1x0+9.1.0 -10.0.-1\] \[f''(x)=12x^2-84x+142\] set that to zero and they should define your interval of concavity

OpenStudy (amistre64):

might be able to factor out a few things to get the numbers down too

OpenStudy (amistre64):

6,24,72 3,12,35 f'(x) = 3x^2 -12x +35